susan invests twice as much money at 7.5% as she does at 6%. If her total interest after a year is $840,how much does she have invested at each rate?
Answer by mananth(16948)

susan invests twice as much money at 7.5% as she does at 6%. If her total interest after a year is $840,how much does she have invested at each rate?
...
6% ------------- x
7.5%-------------2x
..
interest = 840
..
to simplify multiply the equation by 100
6x+7.5*2x=840*100
6x+15x=84000
21x=84000
/21
x=84000/21
x=$4000 at 6%
$8000 at7.5 %
m.ananth@hotmail.ca